Total 39 1007.661 Coefficients Standard Error t Stat р.value Intercept 45.768 3.861 11.854 0.000 Displacement 0.0176 0.00883 xxx 0.054 Weight 0.00361 0.00074 x 0.000 Automatic trans. -0.654 1.201 -0.5450.589 The researcher wants to predict the miles/gallon for an automobile that has automatic transmission, displacement is 250 (cubic in.) and weight is 3000 pound. Based on the regression output, what is the predicted miles/gallon? 35.40 21.03 29.88 34.72 O 30.59
Answer #1

Correct option:

29.88

EXPLANATION:

The Regression Equation is:

Miles = 45.768-0.0176Displacementー0.00361Weǐqht_0.654Automatictransmission

Given:

Displacement = 250

Weight = 3000

Automatic transmission = 1

Substituting, we get:

Miles = 45.768-(0.0176 × 250)-(0.00361 × 3000)-(0.654 × 1)

45.768-4.4-10.83 -0.654

= 45.768 - 15.884

= 29.88